Lisa is weighing combinations of children's blocks. She forgot to write down the weights but thinks that she remembers them correctly. Four blue blocks and five red blocks weigh 17 ounces. One blue block and eight red blocks weigh 38 ounces. What is the combined weight of one red block and one blue block?



1
Expert's answer
2021-02-01T07:13:02-0500

4 blue blocks + 5 red blocks = 17 ounces


4B+ 5R = 17 (equation 1)


1 blue block + 8 red blocks = 38 ounces


1B + 8R = 38 (equation 2)


Now multiply 'equation 2' by 4 and subtract 'equation1' from 'equation 2'. Thus we get 

(4B+32R) - (4B+5R) = (38×4) - (17)


27R = 135, R = 5

Put R = 5 in equation 2, then we get


B+ 40 = 38, B = -2

Thus, weight of combined one blue block and one red block is

1B + 1R = -2 + 5 = 3